OpenStudy (aaronandyson):

-2 le \frac{ 1 }{ 2 } - \frac{ 2x }{ 3 } < 1 \frac{ 5 }{ 6 }

OpenStudy (aaronandyson):

-2 le \frac{ 1 }{ 2 } - \frac{ 2x }{ 3 } < 1 \frac{ 5 }{ 6 }

OpenStudy (aaronandyson):

\[-2 \le \frac{ 1 }{ 2 } - \frac{ 2x }{ 3 } < 1 \frac{ 5 }{ 6 }\]

OpenStudy (aaronandyson):

@Kainui

OpenStudy (aaronandyson):

@ganeshie8

OpenStudy (aaronandyson):

@iambatman

OpenStudy (anonymous):

Hey, have you tried anything? In this case you just have to do what ever you doing in the middle to both sides, so add 1/2 to both sides to get rid of it from the middle.

ganeshie8 (ganeshie8):

maybe start by multiplying \(2*3\) through out so that the fractions go away

ganeshie8 (ganeshie8):

\[-2 \le \frac{ 1 }{ 2 } - \frac{ 2x }{ 3 } < 1 \frac{ 5 }{ 6 }\] \[-2(\color{blue}{2*3}) \le (\color{blue}{2*3})\frac{ 1 }{ 2 } - (\color{blue}{2*3})\frac{ 2x }{ 3 } < (\color{blue}{2*3})1 \frac{ 5 }{ 6 }\]

OpenStudy (anonymous):

\[-2 \le \frac{ 1 }{ 2 } - \frac{ 2x }{ 3 } < \frac{11 }{ 6 }\]\[-2 -\frac{ 1 }{ 2 }\le \frac{ 1 }{ 2 } - \frac{ 2x }{ 3 }-\frac{ 1 }{ 2 } < \frac{11 }{ 6 }-\frac{ 1 }{ 2 }\]\[-\frac{ 3 }{ 2 } \le -\frac{ 2x }{3 }<\frac{ 8 }{ 6 }\]\[-9 \le-4x <8\]now divide all sides by -4 Remember to flip the inequality signs when you divide by a negative number. a<b will become -a>-b.